OpenStudy (anonymous):

a)Construct a function f(x) that is only differentiate at x =0 and x =1 b) More generally, given a1,.....,an in R distinct. Construct a function g(x) that is differentiable at x = a1, a2, ...., an Please, help

OpenStudy (anonymous):

@perl

OpenStudy (idku):

you want a function, e.g. that will be continuous over [0,1] but not any farther? I would think if you just define a function in a couple of ways. Make up/give it a definition you feel like. (i might be wrong)

OpenStudy (anonymous):

@idu its not differential on an interval see distinsit points 0 and 1 right @OOOPS ?

OpenStudy (anonymous):

@idu a function differentiate at a, then it is continuous at a but not vice versa.

OpenStudy (idku):

sure, tnx for showing:)

OpenStudy (anonymous):

@Marki yes, they are distinct points

OpenStudy (anonymous):

eh, i need ur help a bit , what topic is it ?

OpenStudy (anonymous):

it is called "Diffentiability of function of one variable"

OpenStudy (anonymous):



OpenStudy (anonymous):

ok wait i'll read about it , i guess i forgot this thing

OpenStudy (anonymous):

My prof uses limit approaching.

OpenStudy (anonymous):

eh seems i dont understand the question it self :| straggling

OpenStudy (anonymous):

ugh lol ok i misunderstood it need not to be discrete xD so we can make it continues function

OpenStudy (rational):

take dirichlet function function and multiply it by x^2(x-1)^2

OpenStudy (anonymous):

show me, please

OpenStudy (rational):

dirichlet function(D(x)) is not continuous anywhere, just show that f(x) = D(x)*x^2(x-1)^2 is differentiable at x = 0 and x = 1 using the limit definition of derivative

OpenStudy (anonymous):

@rational so x^2(x-1)^2 cuz of (x-0)^2(x-1)^2?

OpenStudy (anonymous):

@rational thanks for the link.

OpenStudy (anonymous):

interesting

OpenStudy (anonymous):

Wonder how to graph: \(f(x) =\begin{cases}x^2~~if~~x\in Q\\0~~if~~x\notin Q\end{cases}\)

OpenStudy (anonymous):

feels something like this XD